# Decoupled structure-preserving discretization of incompressible MHD equations with general boundary conditions

### v1

@@ 3.2.3 Weak conservation of Gauss’s law for magnetism
The second and third terms on the left hand side are zero because we strongly satisfy ×$\nabla\times\nabla q_{h}=0$ since our discrete function spaces constitute a finite dimensional de Rham complex. We are left with the following equation,
@@ 3.2.4 Energy Conservation
(17) $\widehat{u}=0\;,\quad\Gamma_{\widehat{\bm{u}}}=\Gamma_{\widehat{\bm{E}}}=\partial\Omega\;,\quad\Gamma_{\widehat{\bm{\omega}}}=\Gamma_{\widehat{\bm{H}}}=\emptyset\;.$
@@
In (21), (i) represents amount of mechanical work done on the system per unit time, (ii) and (iii) are dissipation rates due to viscosity and electric resistance, respectively, and terms (iv) - (vi) are the net flux of energy through the domain boundary.
Clearly, if $\bm{f}=\bm{0}$, when the MHD flow is ideal (i.e. $\mathrm{R}_{f}=\mathrm{R}_{m}=\infty$) and there is no net flux of energy through the domain boundary, formulation (11) preserves energy, i.e.,
@@
4 Fully-discrete weak formulation
If a conservative temporal discrete scheme, for example, the Crank-Nicolson scheme, is applied to the semi-discrete formulation (11) directly, we can obtain a fully-discrete formulation of same conservation and dissipation properties (except that the time derivative of energy is now expressed in a temporally discrete format). The proof is straightforward; it mimics the analyses in Section 3.2. Therefore, the rest of this paper is devoted to a decoupled formulation with computational efficiency in mind.
@@ 4.1 Decoupled temporal discretization
Consider a set of time instants,
$\left\{t^{0},t^{\frac{1}{2}},t^{1},t^{1+\frac{1}{2}},t^{2},\cdots\right\},$
@@
A fully discrete, temporally decoupled formulation of (11) (hereinafter referred to as the decoupled formulation) is expressed as follows. Given $\bm{f}\in\left[L^{2}(\Omega)\right]^{3}$, natural boundary conditions $\widehat{P}\in H^{1/2}\left(\Omega;\Gamma_{\widehat{P}}\right),\ \widehat{\bm{u}}\in\mathcal{T}H\left(\mathrm{curl};\Omega,\Gamma_{\widehat{\bm{u}}}\right),\ \widehat{\bm{E}}\in\mathcal{T}H\left(\mathrm{curl};\Omega,\Gamma_{\widehat{\bm{E}}}\right)$, and initial conditions $\left(\bm{u}_{h}^{0},\bm{\omega}_{h}^{0},\bm{H}^{\frac{1}{2}}_{h}\right)\in D\left(\Omega\right)\times C\left(\Omega\right)\times C\left(\Omega\right)$, for time step index $k=1,2,3,\cdots$ successively,
@@
(step 1) seek $\left(\bm{u}^{k}_{h},\bm{\omega}^{k}_{h},P^{k-\frac{1}{2}}_{h}\right)\in D_{\widehat{u}}(\Omega,\Gamma_{\widehat{u}})\times C_{\widehat{\bm{\omega}}}^{\parallel}(\Omega,\Gamma_{\widehat{\bm{\omega}}})\times S(\Omega)$, such that, $\forall\left(\bm{v}_{h},\bm{w}_{h},q_{h}\right)\in D_{0}\left(\Omega,\Gamma_{\widehat{u}}\right)\times C^{\parallel}_{\bm{0}}\left(\Omega,\Gamma_{\widehat{\bm{\omega}}}\right)\times S\left(\Omega\right)$,

### v2

@@ 3.2.3 Weak conservation of Gauss’s law for magnetism
One of the most common forms of the MHD equations is the conservative form (or divergence form); see, for example, [24]. The discretization of the conservative form poses several challenges. One of these challenges is the involution constraint $\nabla\cdot\bm{H}=0$ (or $\nabla\cdot\bm{B}=0$), which has been widely discussed in the literature, e.g. [8, 10, 25, 26, 27, 28]. The main root of this challenge is that this formulation is only valid if the involution constraint is satisfied [24, 26]. Moreover, in conservation form, the involution constraint is not an implication of these equations and, therefore, must be enforced. On the other hand, the formulation we employ in this work is self-consistent with regard to the involution constraint. By self-consistent, we mean the following: The involution constraint must only be imposed on the initial condition. Once this condition is satisfied by the initial condition, the evolution equation enforces it at all time instants. As long as numerical discretization preserves a discrete de Rham complex, this equation will automatically be satisfied and the involution constraint will be enforced if the initial condition is divergence-free; see (4).
In a different context, it has been reported that the formulation leads to better results than the conservation form [26], when not enforcing the involution constraint. More recently, [24], a similar weak formulation has been proposed. In that work, as in this, the weak divergence-free constraint is preserved throughout the whole evolution. As seen in the next section (and in the numerical results), this does not lead to a loss of energy conservation, in contrast to approaches for the conservative form that do not strongly satisfy the divergence-free constraint.
@@
The second and third terms on the left-hand side are zero because we strongly satisfy ×$\nabla\times\nabla q_{h}=0$ since our discrete function spaces constitute a finite-dimensional de Rham complex. We are left with the following equation,
@@ 3.2.4 Energy conservation
(17) $\hat{u}=0\;,\quad\Gamma_{\widehat{\bm{u}}}=\Gamma_{\widehat{\bm{E}}}=\partial\Omega\;,\quad\Gamma_{\widehat{\bm{\omega}}}=\Gamma_{\widehat{\bm{H}}}=\emptyset\;.$
@@
In (21), (i) represents the amount of mechanical work done on the system per unit time, (ii) and (iii) are dissipation rates due to viscosity and electric resistance, respectively, and terms (iv) - (vi) are the net flux of energy through the domain boundary. Clearly, if $\bm{f}=\bm{0}$, when the MHD flow is ideal (i.e. $\mathrm{R}_{f}=\mathrm{R}_{m}=\infty$) and there is no net flux of energy through the domain boundary, formulation (11) preserves energy, i.e.
@@
4 Fully discrete weak formulation
If a conservative temporal discrete scheme, for example, the Crank-Nicolson scheme, is applied to the semi-discrete formulation (11) directly, we can obtain a fully discrete formulation of the same conservation and dissipation properties (except that the time derivative of energy is now expressed in a temporally discrete format). See B. The proof is straightforward; it mimics the analyses in Section 3.2. Therefore, the rest of this paper is devoted to a decoupled formulation with computational efficiency in mind.
@@ 4.1 Decoupled temporal discretization
Consider a set of time instants
(22) $\left\{t^{0},t^{\frac{1}{2}},t^{1},t^{1+\frac{1}{2}},t^{2},\cdots\right\},$
@@
A fully discrete, temporally decoupled formulation of (11) (hereinafter referred to as the decoupled formulation) is expressed as follows. Given $\bm{f}\in\left[L^{2}(\Omega)\right]^{3}$, natural boundary conditions $\widehat{P}\in H^{1/2}\left(\Omega;\Gamma_{\widehat{P}}\right),\ \widehat{\bm{u}}\in\mathcal{T}H\left(\mathrm{curl};\Omega,\Gamma_{\widehat{\bm{u}}}\right),\ \widehat{\bm{E}}\in\mathcal{T}H\left(\mathrm{curl};\Omega,\Gamma_{\widehat{\bm{E}}}\right)$, and initial conditions $\left(\bm{u}_{h}^{0},\bm{\omega}_{h}^{0},\bm{H}^{\frac{1}{2}}_{h}\right)\in D\left(\Omega\right)\times C\left(\Omega\right)\times C\left(\Omega\right)$, for time-step index $k=1,2,3,\cdots$ successively,
@@
(step 1) seek $\left(\bm{u}^{k}_{h},\bm{\omega}^{k}_{h},P^{k-\frac{1}{2}}_{h}\right)\in D_{\hat{u}}(\Omega,\Gamma_{\hat{u}})\times C_{\widehat{\bm{\omega}}}^{\parallel}(\Omega,\Gamma_{\widehat{\bm{\omega}}})\times S(\Omega)$, such that, $\forall\left(\bm{v}_{h},\bm{w}_{h},q_{h}\right)\in D_{0}\left(\Omega,\Gamma_{\hat{u}}\right)\times C^{\parallel}_{\bm{0}}\left(\Omega,\Gamma_{\widehat{\bm{\omega}}}\right)\times S\left(\Omega\right)$,